# Mixed Dynamics In Linear Networks: Unifying the Lazy and Active Regimes

### v1

@@ A.3 Perturbation of Singular Values and Singular Vectors
Appendix B Gradient Flow Dynamics of $A_{t}$ in Active Regime
B.1 Saddle to Saddle Regime
@@
The goal of this section is to prove that the singular vectors of $A_{t}$ is well-aligned with the singular vectors of $A^{*}$, throughout the Saddle-to-Saddle regime. In the rest of the section, we will assume the dependence of $A_{t}$ on $t$ and use $A$ to represent $A_{t}$.
Definition B.1.
@@
A simple result of the induction lemma B.3, lemma B.4, lemma B.6 and lemma B.7. ∎
@@
B.2 First NTK Regime
At initialization, the singular values of $A$ are of order $\sigma^{2}\sqrt{wd}$ by using Mechenko-Pastur law on $A^{T}A$, which is infinitely smaller than $\sigma^{2}w$. As a result, there is a very short period when the dynamics is very close to NTK.
Lemma B.8.
@@ B.2 First NTK Regime
By lemma B.4 we have $x\leq Cd^{-\text{min}(\frac{1}{2},\frac{\delta}{2})}$ for some constant $C$. ∎
Lemma B.9.
@@
Let $t=t_{2}$ we see that |$\absolutevalue{\|A(t_{2})\|_{op}-\|B(t_{2})\|_{op}}\leq\|A-B\|_{op}(t_{2})\leq 2\frac{\|A_{t_{2}}\|^{2}_{op}}{\sigma^{2}w}a_{1}c$. Therefore as long as $c$ is sufficiently small, we can guarantee that $\|A-B\|_{op}(t)\leq\frac{\text{min}_{1\leq k\leq K}(a_{k-1}-a_{k})}{100a_{1}}\|B\|_{op}(t)$. This guarantees that the singular values of $A$ grows linearly in $[t_{1},t_{2}]$, that $s_{k}-s_{k+1}\geq cs_{k}$ if $a_{k}>a_{k+1}$ for some constant $c$, that $s_{K+1}<\frac{1}{2}s_{K}$ and that $A_{t_{2}}\in P_{1}$. It remains to check that $A_{t_{2}}$ satisfies $P_{2}(C,\frac{\gamma_{w}-1}{2})$. We are ready to use similar techniques as in lemma B.6. All computation are similar, and the only difference is that |$\absolutevalue{R(k,j)}=O(\frac{\sigma^{2}w}{s_{K}})$. We have
@@
Therefore $A_{[t_{1},t_{2}]}\subset P_{2}(C,\text{min}(\frac{\gamma_{w}-1}{4},\frac{1}{2}))$. Let $\gamma=\text{min}(\frac{\gamma_{w}-1}{4},\frac{1}{2}).$ From lemma B.8 we see that if $n_{k_{1}}+1\leq p_{1},p_{2}\leq n_{k}$, then |$\absolutevalue{s_{p_{1}}-s_{p_{2}}}\leq O(d^{-\gamma})s_{K}$. Moreover, the dynamics of $s_{p_{i}},i=1,2$ are both given by
@@
Appendix C Proof of Theorem 1
C.1 Weak bound
We prove the following weak bound.
Proposition C.1.
For every $\varepsilon>0$ and every $t<T$, we have with high probability,
Analogous results holds for $W_{2}$.
Our main tool is the following lemma.
Lemma C.2.
For every cost $C$ and every time $t$, we have
$W_{2}^{T}W_{2}(t)-W_{1}W_{1}^{T}(t)=W_{2}^{T}(0)W_{2}(0)-W_{1}(0)W_{1}^{T}(0)$ (10)
Proof.
Let $L=\|A^{*}-A\|^{2}$ be the loss function. Then
$\partial_{t}W_{1}$ $=2W_{2}^{T}\nabla C$
$\partial_{t}W_{2}$ $=2\nabla CW_{1}^{T}$
We see that
Lemma C.3.
  1. 1.
    The image of $P_{1}$ and $P_{2}$ are orthogonal to each other;
  2. 2.
    With high probability,
Proof of Lemma C.3 .
  1. 1.
    $\mathbb{E}[X_{i\mu}]=0$;
  2. 2.
    $\mathbb{E}[\absolutevalue{X_{i\mu}}^{2}]=\frac{1}{\sqrt{MN}}$;
  3. 3.
    for every $p\in\mathbb{N}$, there exists a constant $C_{p}$ such that
Here, $M$ satisfies
$0<C^{-1}\leq\frac{\log M}{\log N}\leq C<\infty$
where
$\gamma_{\pm}:=\sqrt{\phi}+\frac{1}{\sqrt{\phi}}\pm 2$ (13)
with high probability. Here $\gamma_{\alpha}$ is defined through
$\int_{\gamma_{k}}^{\infty}\rho_{\phi}(dx)=\frac{k}{d}$ (15)
[…]
$\leq O(\sigma^{2}\sqrt{wd})$ (17)
Similarly, there exists a projection $\tilde{P}_{2}$ such that
$v_{k}^{\prime}=v_{k}-P_{1}v_{k}.$ (20)
We conclude that for each $\ell$ we have, by CLT,
In particular,
$\leq$ $2w\mathbb{P}\{N(0,1)>100\log d\}$
$\leq$ $O(d^{-50})$
$\|P_{1}v_{k}\|\leq\sigma\sqrt{d}\log d$ (26)
Proof of Proposition C.1.
: We have
$=A^{T}A+W_{1}^{T}(W_{1}W_{1}^{T}(0)-W_{2}^{T}W_{2}(0))W_{1}$
From lemma C.3 we see that as positive semi-definite matrix, for every constant $\varepsilon>0$,
$0\leq W_{1}W_{1}^{T}(0)\leq(1+\varepsilon)\sigma^{2}wI$ (27)
with high probability. Therefore
By moving terms around and corollary, we have
$\leq$ $A^{T}A+\frac{(1+\varepsilon)^{2}}{4}\sigma^{4}w^{2}I$ (30)
C.2 Strong Bound
Recall that the evolution of weight matrix in gradient descent is given by the following.
$\frac{d}{dt}W_{1}(t)=\eta W_{2}^{T}\nabla C$ (33)
$\frac{d}{dt}W_{2}(t)=\eta\nabla CW_{1}^{T}$ (34)
The goal of this section is to prove that
$W_{1}^{T}W_{1}\approx\sqrt{A^{T}A+\sigma^{4}w^{2}I}$ (35)
Proposition C.4.
For every cost $C$ we have
Proof of Lemma C.4 .
: We start from the equations:
Recall that $\frac{dw}{dt}=$ Substracting the second pair of equations from the first pair and denoting $dC_{i}=C_{i}-\hat{C}$, we obtain:
$\|C_{1}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=C_{1}^{3}-\hat{C}_{1}^{3}-A^{T}AdC_{1}-dC_{1}A^{T}A-\sigma^{4}w^{2}dC_{1}+A^{T}dC_{2}A$
$\|C_{2}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=C_{2}^{3}-\hat{C}_{2}^{3}-AA^{T}dC_{2}-dC_{2}AA^{T}-\sigma^{4}w^{2}dC_{2}+AdC_{1}A^{T}.$
Now since
$C_{1}^{3}-\hat{C}_{1}^{3}$ $=\hat{C}_{1}^{2}dC_{1}+\hat{C}_{1}dC_{1}C_{1}+dC_{1}C_{1}^{2}$
we obtain
$\|C_{1}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=\left(\hat{C}_{1}^{2}-A^{T}A\right)dC_{1}+\hat{C}_{1}dC_{1}C_{1}+dC_{1}\left(C_{1}^{2}-A^{T}A\right)-dC_{1}+A^{T}dC_{2}A$
$=\hat{C}_{1}dC_{1}C_{1}+dC_{1}\left(C_{1}^{2}-A^{T}A\right)+A^{T}dC_{2}A,$
and similarly for the other index. For any singular value $s_{i}$ of $A$, with left and right singular vectors $u_{i},v_{i}$, we multiply the first equation to the left by $v_{i}^{T}$ and the second by $u_{i}^{T}$ to obtain an equation for $v_{i}^{T}dC_{1}$ and $u_{i}^{T}dC_{2}$:
To show that $\|v_{i}^{T}dC_{1}\|$ and $\|u_{i}^{T}dC_{2}\|$, we need to show that the $(d_{in}+d_{out})\times(d_{in}+d_{out})$ block matrix
is full rank. This matrix can be further simplified to
The inner matrix can then be rewritten as $RR^{T}$ where $R$ is defined as
Let $Q=\left(\begin{array}[]{cc}W_{1}^{T}&0\\ 0&W_{2}\end{array}\right)$. Then the smallest nonzero singular value of $QR(QR)^{T}$ is the same as the smallest nonzero singular value of $(QR)^{T}QR$. We expand $(QR)^{T}QR$ as follows.
$(QR)^{T}QR$
where we used the fact that
Tthe $(d_{in}+d_{out})$-th eigenvalue of the above is lower bounded by $\sigma^{2}w\sqrt{\left(\frac{s_{i}}{\sigma^{2}w}\right)^{2}+1}+\sigma^{2}w-\sigma^{2}w\left(\frac{s_{i}}{\sigma^{2}w}\right)\geq\sigma^{2}w$. We conclude that
$\|dC_{1}\|_{op}+\|dC_{2}\|_{op}\leq O(\sqrt{\frac{d}{w}})\|C_{1}\|_{op},$
Compared to lemma C.3, lemma C.4 gives a tighter bound on $\|C_{1}-\hat{C}_{1}\|_{\mathrm{op}}$ when $\|C_{1}\|_{\mathrm{op}}\leq\sigma^{2}w\sqrt{\frac{d}{w}}$.

### v2

@@ A.3 Perturbation of Singular Values and Singular Vectors
Appendix B Proof of Theorem 1
B.1 Weak bound
We prove the following weak bound.
Proposition B.1.
For every $\varepsilon>0$ and every $t<T$, we have with high probability,
Analogous results holds for $W_{2}$.
Our main tool is the following lemma.
Lemma B.2.
For every cost $C$ and every time $t$, we have
$W_{2}^{T}W_{2}(t)-W_{1}W_{1}^{T}(t)=W_{2}^{T}(0)W_{2}(0)-W_{1}(0)W_{1}^{T}(0)$ (4)
Proof.
Let $L=\|A^{*}-A\|^{2}$ be the loss function. Then
$\partial_{t}W_{1}$ $=2W_{2}^{T}\nabla C$
$\partial_{t}W_{2}$ $=2\nabla CW_{1}^{T}$
We see that
Lemma B.3.
  1. 1.
    The image of $P_{1}$ and $P_{2}$ are orthogonal to each other;
  2. 2.
    With high probability,
Proof of Lemma B.3 .
  1. 1.
    $\mathbb{E}[X_{i\mu}]=0$;
  2. 2.
    $\mathbb{E}[\absolutevalue{X_{i\mu}}^{2}]=\frac{1}{\sqrt{MN}}$;
  3. 3.
    for every $p\in\mathbb{N}$, there exists a constant $C_{p}$ such that
Here, $M$ satisfies
$0<C^{-1}\leq\frac{\log M}{\log N}\leq C<\infty$
where
$\gamma_{\pm}:=\sqrt{\phi}+\frac{1}{\sqrt{\phi}}\pm 2$ (7)
with high probability. Here $\gamma_{\alpha}$ is defined through
$\int_{\gamma_{k}}^{\infty}\rho_{\phi}(dx)=\frac{k}{d}$ (9)
[…]
$\leq O(\sigma^{2}\sqrt{wd})$ (11)
Similarly, there exists a projection $\tilde{P}_{2}$ such that
$v_{k}^{\prime}=v_{k}-P_{1}v_{k}.$ (14)
We conclude that for each $\ell$ we have, by CLT,
In particular,
$\leq$ $2w\mathbb{P}\{N(0,1)>100\log d\}$
$\leq$ $O(d^{-50})$
$\|P_{1}v_{k}\|\leq\sigma\sqrt{d}\log d$ (20)
Proof of Proposition B.1.
: We have
$=A^{T}A+W_{1}^{T}(W_{1}W_{1}^{T}(0)-W_{2}^{T}W_{2}(0))W_{1}$
From lemma B.3 we see that as positive semi-definite matrix, for every constant $\varepsilon>0$,
$0\leq W_{1}W_{1}^{T}(0)\leq(1+\varepsilon)\sigma^{2}wI$ (21)
with high probability. Therefore
By moving terms around and corollary, we have
$\leq$ $A^{T}A+\frac{(1+\varepsilon)^{2}}{4}\sigma^{4}w^{2}I$ (24)
B.2 Strong Bound
The weak bound does not provide useful information if $\|W_{1}^{T}W_{1}\|_{op}<<\sigma^{2}w$. For this reason we prove strong bound, which provide useful information if $\|W_{1}^{T}W_{1}\|_{op}<<\sigma^{2}w^{1+\Box}$ for some constant $\Box>0$. Recall that the evolution of weight matrix in gradient descent is given by the following.
$\frac{d}{dt}W_{1}(t)=\eta W_{2}^{T}\nabla C$ (27)
$\frac{d}{dt}W_{2}(t)=\eta\nabla CW_{1}^{T}$ (28)
The goal of this section is to prove that
$W_{1}^{T}W_{1}\approx\sqrt{A^{T}A+\sigma^{4}w^{2}I}$ (29)
Proposition B.4.
For every cost $C$ we have
Proof of Lemma B.4 .
: We start from the equations:
Apriori, the cubic equation for $C_{1}$ and $C_{2}$ might have multiple solutions. We give an intuitive argument explaining why $\hat{C}_{1}$ and $\hat{C}_{2}$ are the correct solutions. By selecting a proper basis, we assume $A=diag(a_{1},\ldots,a_{d})$ is diagonal. Assume that $(W_{2}^{T}W_{2}-W_{1}W_{1}^{T})^{2}=\sigma^{2}wI$. Also assume that $C_{1}$ and $C_{2}$ both commute with $A$. In this case the equations for $C_{1}$ and $C_{2}$ reduces to cubic equation for scalars. Let $\lambda_{1},\ldots,\lambda_{d}$ be the eigenvalues of $C_{1}$. Solving the equations for scalars, we have $\lambda_{i}=0$ or ±$\pm\sqrt{a_{i}^{2}+\sigma^{4}w^{2}}$. By lemma B.2 and lemma B.3, we have
$W_{1}W_{1}^{T}(t)-W_{2}^{T}W_{2}(t)=W_{1}W_{1}^{T}(0)-W_{2}^{T}W_{2}(0)=\sigma^{2}wP_{1}-\sigma^{2}wP_{2}+o(\sigma^{2}w).$
Since $W_{1}W_{1}^{T}$ is positive semi-definite, all its eigenvalues are non-negative, and thus
$W_{1}W_{1}^{T}(t)\geq(\sigma^{2}w+o(\sigma^{2}w))P_{1}.$
Since the top $w$ eigenvalues of $W_{1}^{T}W_{1}$ is the same as the top $w$ eigenvalues of $W_{1}W_{1}^{T}$, we conclude that $\lambda_{i}\geq\sigma^{2}w(1+o(1))$. This forces $\lambda_{i}=\sqrt{a_{i}^{2}+\sigma^{4}w^{2}}$.
Since $C_{1}$, $C_{2}$ are not assumed to be aligned with $A$, we cannot reduce the matrix cubic equations into scalar cubic equations. The high level idea for proving $dC_{i}:=C_{i}-\hat{C}_{i}$ is small is the inverse function theorem.
  1. 1.
    Step 1: show that LHS of the equations are small.
  2. 2.
    Step 2: reduce the RHS of the equations to a linear function of $dC_{1}$ and $dC_{2}$. The system of equations is thus reduced to
    $\begin{pmatrix}\text{small}\\ \text{small}\end{pmatrix}=\begin{pmatrix}*&*\\ *&*\end{pmatrix}\begin{pmatrix}v_{i}^{T}dC_{1}\\ u_{i}^{T}dC_{2}\end{pmatrix}.$
    The * matrix is now the "Jacobian" matrix, and $u_{i},v_{i}$ are left and right singular vectors of $A$.
  3. 3.
    Step 3: prove that the "Jacobian" matrix ($\ *&*\end{pmatrix}$ is strictly positive definite, thus proving that $v_{i}^{T}dC_{1}$ and $u_{i}^{T}dC_{2}$ have small magnitude for all $i$.
Substracting the second pair of equations from the first pair and denoting $dC_{i}=C_{i}-\hat{C}$, we obtain:
$\|C_{1}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=C_{1}^{3}-\hat{C}_{1}^{3}-A^{T}AdC_{1}-dC_{1}A^{T}A-\sigma^{4}w^{2}dC_{1}+A^{T}dC_{2}A$ (31)
$\|C_{2}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=C_{2}^{3}-\hat{C}_{2}^{3}-AA^{T}dC_{2}-dC_{2}AA^{T}-\sigma^{4}w^{2}dC_{2}+AdC_{1}A^{T}.$ (32)
Now since
$C_{1}^{3}-\hat{C}_{1}^{3}$ $=\hat{C}_{1}^{2}dC_{1}+\hat{C}_{1}dC_{1}C_{1}+dC_{1}C_{1}^{2},$
we substitute the above relation to equation 31 and obtain
$\|C_{1}\|_{op}O(\sigma^{4}w^{2}\sqrt{\frac{d}{w}})$ $=\left(\hat{C}_{1}^{2}-A^{T}A\right)dC_{1}+\hat{C}_{1}dC_{1}C_{1}+dC_{1}\left(C_{1}^{2}-A^{T}A\right)-dC_{1}+A^{T}dC_{2}A$ (33)
$=\hat{C}_{1}dC_{1}C_{1}+dC_{1}\left(C_{1}^{2}-A^{T}A\right)+A^{T}dC_{2}A,$ (34)
and similarly for equation 32. For any singular value $s_{i}$ of $A$, with left and right singular vectors $u_{i},v_{i}$, we multiply equation 34 to the left by $v_{i}^{T}$, and divide both sides by $\sigma^{2}w$, to obtain an equation for $v_{i}^{T}dC_{1}$. Similarly, we obtain an equation for $u_{i}^{T}dC_{2}$: for $v_{i}^{T}dC_{1}$ and $u_{i}^{T}dC_{2}$:
To show that $\|v_{i}^{T}dC_{1}\|$ and $\|u_{i}^{T}dC_{2}\|$ are small, it suffices to show that the $(d_{in}+d_{out})\times(d_{in}+d_{out})$ block matrix
is strictly positive definite. This matrix can be further simplified to
The inner matrix can then be rewritten as $RR^{T}$ where $R$ is defined as
Let $Q=\left(\begin{array}[]{cc}W_{1}^{T}&0\\ 0&W_{2}\end{array}\right)$.The "Jacobian" matrix is then $QR(QR)^{T}$. As described in the strategy, we need to show that the singular values are strictly positive. The smallest nonzero singular value of $QR(QR)^{T}$ is the same as the smallest nonzero singular value of $(QR)^{T}QR$. We expand $(QR)^{T}QR$ as follows.
$(QR)^{T}QR$
where we used the fact that
The $(d_{in}+d_{out})$-th eigenvalue of the above is lower bounded by $\sigma^{2}w\sqrt{\left(\frac{s_{i}}{\sigma^{2}w}\right)^{2}+1}+\sigma^{2}w-\sigma^{2}w\left(\frac{s_{i}}{\sigma^{2}w}\right)\geq\sigma^{2}w$. We conclude that
$\|dC_{1}\|_{op}+\|dC_{2}\|_{op}\leq O(\sqrt{\frac{d}{w}})\|C_{1}\|_{op},$
Compared to lemma B.3, lemma B.4 gives a tighter bound on $\|C_{1}-\hat{C}_{1}\|_{\mathrm{op}}$ when $\|C_{1}\|_{\mathrm{op}}\leq\sigma^{2}w\sqrt{\frac{d}{w}}$.
As suggested by an anonymous referee, it is possible to obtain the same approximated dynamics of $A_{\theta(t)}$ by imposing a non-homogeneous balance condition (in a different setup). Assume that $W_{1}W_{1}^{T}-W_{2}^{T}W_{2}=2\sigma^{2}wI$. Then
$C_{1}^{2}+2\sigma^{2}wC_{1}-A^{T}A=0;$
$C_{2}^{2}-2\sigma^{2}wC_{2}-AA^{T}=0.$
Therefore $C_{1}=-\sigma^{2}w+\sqrt{A^{T}A+\sigma^{4}w^{2}}$ and $C_{2}=\sigma^{2}w+\sqrt{AA^{T}+\sigma^{4}w^{2}}$. Substituting into Gradient Flow equation, we have
$\frac{dA}{dt}=-\eta(\sqrt{AA^{T}+\sigma^{4}w^{2}}\nabla C+\nabla C\sqrt{A^{T}A+\sigma^{4}w^{2}}).$
The advantage of the setup is that it significantly simplifies the proof. The The limitation of the setup is that $W_{1}W_{1}^{T}-W_{2}^{T}W_{2}\neq\sigma^{4}w^{2}I$ if the initial variance of entries of $W_{1}$ and $W_{2}$ are comparable. In this case, $W_{1}W_{1}^{T}-W_{2}^{T}W_{2}$ will have $w$ positive singular values and $w$ negative singular values, and the absolute value of positive and negative singular values are comparable. In the setup of our problem, the variance of entries of $W_{1}$ and $W_{2}$ equal.
Appendix C Gradient Flow Dynamics of $A_{t}$ in Active Regime
C.1 Saddle to Saddle Regime
@@
The goal of this section is to prove that the singular vectors of $A_{t}$ is well-aligned with the singular vectors of $A^{*}$, throughout the Saddle-to-Saddle regime. In the rest of the section, we will assume the dependence of $A_{t}$ on $t$ and use $A$ to represent $A_{t}$. If at initialization, $A_{t}$ commutes with $A^{*}$, then throughout the training, $A_{t}$ will always commute with $A^{*}$. In this section, we use a delicate stability argument to show that if $A_{t}$ almost commute with $A^{*}$ at the beginning of the Saddle to Saddle regime, then it will continue to be almost commutative with $A^{*}$ throughout the training process.
Definition C.1.
@@
A simple result of the induction lemma C.3, lemma C.4, lemma C.6 and lemma C.7. ∎
@@
C.2 First NTK Regime
At initialization, the singular values of $A$ are of order $\sigma^{2}\sqrt{wd}$ by using Mechenko-Pastur law on $A^{T}A$, which is infinitely smaller than $\sigma^{2}w$. As a result, there is a very short period when the dynamics is very close to NTK. This section is again a delicate stability argument to show that at the end of the first NTK regime, $A_{t}$ is almost commutative with $A^{*}$.
Lemma C.8.
@@ C.2 First NTK Regime
By lemma C.4 we have $x\leq Cd^{-\text{min}(\frac{1}{2},\frac{\delta}{2})}$ for some constant $C$. ∎
Lemma C.9.
@@
Let $t=t_{2}$ we see that |$\absolutevalue{\|A(t_{2})\|_{op}-\|B(t_{2})\|_{op}}\leq\|A-B\|_{op}(t_{2})\leq 2\frac{\|A_{t_{2}}\|^{2}_{op}}{\sigma^{2}w}a_{1}c$. Therefore as long as $c$ is sufficiently small, we can guarantee that $\|A-B\|_{op}(t)\leq\frac{\text{min}_{1\leq k\leq K}(a_{k-1}-a_{k})}{100a_{1}}\|B\|_{op}(t)$. This guarantees that the singular values of $A$ grows linearly in $[t_{1},t_{2}]$, that $s_{k}-s_{k+1}\geq cs_{k}$ if $a_{k}>a_{k+1}$ for some constant $c$, that $s_{K+1}<\frac{1}{2}s_{K}$ and that $A_{t_{2}}\in P_{1}$. It remains to check that $A_{t_{2}}$ satisfies $P_{2}(C,\frac{\gamma_{w}-1}{2})$. We are ready to use similar techniques as in lemma C.6. All computation are similar, and the only difference is that |$\absolutevalue{R(k,j)}=O(\frac{\sigma^{2}w}{s_{K}})$. We have
@@
Therefore $A_{[t_{1},t_{2}]}\subset P_{2}(C,\text{min}(\frac{\gamma_{w}-1}{4},\frac{1}{2}))$. Let $\gamma=\text{min}(\frac{\gamma_{w}-1}{4},\frac{1}{2}).$ From lemma C.8 we see that if $n_{k_{1}}+1\leq p_{1},p_{2}\leq n_{k}$, then |$\absolutevalue{s_{p_{1}}-s_{p_{2}}}\leq O(d^{-\gamma})s_{K}$. Moreover, the dynamics of $s_{p_{i}},i=1,2$ are both given by